OpenStudy (anonymous):

solve the system of equation by graphing x-2y=8 x+2y=-4

OpenStudy (anonymous):

first you'd want to put it in y=mx+b form: x-2y=8 -> 2y=x-8 -> y=(1/2)x-4 ----> slope=(1/2), y-intercept=-4 x+2y=-4 -> 2y=-x-4 -> y=(-1/2)x-2 ----> slope=(-1/2), y-intercept=-2 Using that information you should be able to graph the two lines and see where they intersect to find the solution

OpenStudy (anonymous):